# Distributed Sign Momentum with Local Steps for Training Transformers

### v1

@@ Abstract
Pre-training Transformer models is resource-intensive, and recent studies have shown that sign momentum is an efficient technique for training large-scale deep learning models, particularly Transformers. However, its application in distributed training or federated learning remains underexplored. This paper investigates a novel communication-efficient distributed sign momentum method with local updates. Our proposed method allows for a broad class of base optimizers for local updates, and uses sign momentum in global updates, where momentum is generated from differences accumulated during local steps. We evaluate our method on the pre-training of various GPT-2 models, and the empirical results show significant improvement compared to other distributed methods with local updates. Furthermore, by approximating the sign operator with a randomized version that acts as a continuous analog in expectation, we present an $O(1/\sqrt{T})$ convergence for one instance of the proposed method for nonconvex smooth functions.
@@ 1 Introduction
$\mathop{\mathrm{minimize}}_{\bm{x}\in\mathbb{R}^{d}}f(\bm{x}):=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\bm{\xi}^{(i)}\sim\mathcal{D}_{i}}F_{i}(\bm{x},\bm{\xi}^{(i)}),$
@@
where $F_{i}$ represents the supervised learning model, $\bm{x}$ denotes the model parameters, and $\bm{\xi}^{(i)}$ refers to the data samples drawn from data distribution $\mathcal{D}_{i}$ on worker $i\in[n]$. The problem formulation (1) serves as a natural abstraction for distributed training (Lian et al., 2017) and federated learning (McMahan et al., 2017). A common approach to solving (1) involves workers first computing mini-batch gradients of $F_{i}$, followed by an all-reduce step to aggregate the local mini-batch gradients and obtain the global average. A global optimization step, such as Stochastic Gradient Descent (SGD) (Robbins and Monro, 1951), Momentum (Polyak, 1964), Nesterov’s Accelerated Gradient (NAG) (Nesterov, 1983), Adam (Kingma, 2014) and its weight decay variant AdamW (Loshchilov, 2017), or more recently Lion (Chen et al., 2024b), is then applied. This method is well suited for distributed learning scenarios where data is spread across multiple workers due to size or privacy constraints, and communication is relatively inexpensive.
@@
However, in distributed training across GPU clusters, synchronized communication during the all-reduce step can become a bottleneck due to stragglers (Dean et al., 2012) or slow inter-node and inter-cluster communication. Several approaches have been proposed to mitigate this communication overhead. Decentralized methods, such as gossiping SGD (Jin et al., 2016) and consensus-based distributed SGD (Jiang et al., 2017), use approximate mini-batch gradient averaging instead of exact averaging to reduce idle time. Local updating methods, like local SGD (Lin et al., 2020; Gorbunov et al., 2021b), local momentum (PR-SGD-Momentum) (Yu et al., 2019) and local Adam (Lu et al., 2023), reduce communication frequency to save on bandwidth. Additionally, some local updating methods further optimize performance by overlapping communication and computation (Wang et al., 2019; Sun et al., 2024), utilizing stale synchronized model parameters for local updates to reduce the waiting time for synchronizing the latest local models. All these approaches, although reduce communication and run faster, add perturbations to the optimization process. These perturbations can lead to decreased performance in terms of training and validation losses under the same amount of computation, especially under heterogeneous local data (Wang et al., 2020).
@@
Several methods have been proposed to enhance the performance of local updating techniques. SCAFFOLD (Karimireddy et al., 2020) introduces a control variate that corrects the local SGD direction using global information, eliminating the dependency on local gradient dissimilarity in the convergence guarantees. One of the most relevant methods, BMUF (Chen and Huo, 2016), combines local updates with a global momentum step to accelerate distributed training, without the need for additional communication, such as transmitting the control variate in SCAFFOLD or the local momentum in PR-SGD-Momentum (Yu et al., 2019). SlowMo (Wang et al., 2019) extends BMUF by providing a general framework that operates on local base optimizers, such as local SGD and SGP (Assran et al., 2019), and periodically applies a global momentum step after running the local optimizer for several iterations.
@@
Beyond local updating methods, momentum has long been recognized as a key technique for improving deep learning training, enhancing both optimization and generalization (Sutskever et al., 2013). More recently, Kunstner et al. (2023) suggest that sign momentum may be the key factor behind Adam’s significant advantage over SGD in training Transformers (Vaswani, 2017). Chen et al. (2024b) introduced the Lion (Evolved Sign Momentum) optimization method, discovered through program search, which outperforms the widely used Adam (Kingma, 2014) in various vision and autoregressive language modeling tasks. Furthermore, Kunstner et al. (2023), through experiments with different batch sizes, suggest that Adam’s significant performance advantage over SGD likely stems from its similarity to sign descent with momentum, particularly in Transformers used for language models.
@@
Inspired by the effectiveness of the sign momentum method in training large neural networks, particularly Transformers, and the success of global momentum steps in local updating methods, a natural question is:
@@
Can we extend the sign momentum method to distributed training of Transformers with local updates?
@@ 1.1 Contributions
In this work, we answer this question affirmatively by proposing a framework for distributed sign momentum with local updates, supported by both empirical evaluations and theoretical justifications. Our contributions are as follows:
@@
    We introduce a new framework for distributed sign momentum with local updates. This framework is versatile, allowing any off-the-shelf base optimizers for local updates, while using differences accumulated from these updates to adjust the momentum in the global sign momentum step. Details are provided in Algorithm 1.
@@
    We evaluate the performance of Algorithm 1 in pre-training GPT-2 models of various sizes, demonstrating consistent improvements over other local updating methods. In the special single worker case, Algorithm 1 and its non-signed variant can even outperform AdamW.
@@
    We provide a general theoretical analysis of the proposed framework by approximating the sign operator with a randomized continuous analog in expectation. When specializing the base optimizer as SGD, we obtain an $O\big{(}1/\sqrt{T}\big{)}$ convergence where $T$ is the total global steps.
@@ 1.2 Related work
Communication compression. In distributed optimization, there has been extensive research on compressing communication between workers to reduce overhead. This includes methods such as 1-bit sign SGD (Seide et al., 2014; Bernstein et al., 2018), gradient difference compression (Gorbunov et al., 2021a), compression with error-feedback (Karimireddy et al., 2019), and compression with error reset (Xie et al., 2020), among others. Although this work focuses on local updating methods without communication compression, several 1-bit compression (sign-based) methods, such as SIGNUM (Bernstein et al., 2019) and Major Vote signSGD (Sun et al., 2023), apply (random) sign operations to local momentum and use majority voting for local signs. In contrast, Algorithm 1 employs full-precision communication during the global synchronization step, with the sign taken after aggregation.
@@
$\bm{x}_{t+1}=\bm{x}_{t}-\eta_{t}\text{sign}(\nabla f(\bm{x},\bm{\xi})),$ (1)
@@
$\begin{split}\bm{m}_{t+1}=\beta\bm{m}_{t}+(1-\beta)\nabla f(\bm{x}_{t},\bm{\xi}),\\ \bm{x}_{t+1}=\bm{x}_{t}-\eta_{t}\text{sign}(\bm{m}_{t+1}),\end{split}$ (2)
@@ 2 Distributed sign momentum
where $\bm{\xi}_{t,k}^{(i)}$ is a sample drawn from $\bm{\xi}^{(i)}$. As in Wang et al. (2019), base optimizers can also involve communication steps, such as SGP (Assran et al., 2019), and momentum steps. We denote the final update direction applied to worker $i$ at iteration $(t,k)$ as ${\bm{d}}_{t,k}^{(i)}$.
@@
Note that (5)-(7) mimics the update rule of Lion (Chen et al., 2024a) (also described in in Section A in Appendix), by treating the differences accumulated from local updates as pseudo-stochastic gradients. In both (5) and (7), we scale $(\bm{x}_{t,0}-\bm{x}_{t,\tau})$ by $\frac{1}{\gamma_{t}}$ to make the momentum buffer independent of the learning rate $\gamma_{t}$, which can vary over time when using a learning rate schedule.
@@
9:     Update global parameter: $\qquad\bm{u}_{t+1}\leftarrow\beta_{1}\bm{m}_{t}+\frac{1-\beta_{1}}{\gamma_{t}}(\bm{x}_{t,0}-\bm{x}_{t,\tau})$ $\qquad\bm{x}_{t+1,0}\leftarrow\bm{x}_{t,0}-\eta\gamma\big{(}\text{sign}(\bm{u}_{t+1})+\lambda\bm{x}_{t,0}\big{)}$
@@
Distributed optimizer. We note that Algorithm 1 can be seamlessly integrated with the ZERO-series distributed optimizer (Rajbhandari et al., 2020) to handle large models. Consider a layered distributed optimization strategy: GPUs on a single node can be treated as a joint worker, utilizing a distributed optimizer like ZERO2 for local updates of the base optimizer, while the global buffers $\bm{x}_{t,0}$ and $\bm{m}_{t,0}$ are distributed across nodes. This approach allows more frequent communication for local updates to leverage faster intra-node communication, such as NVLinks.
Table 1: Model configurations and peak learning rates
Size Embedding dimension # of heads Depth Peak LR
Small 125M 768 12 12 5e-4
Medium 355M 1024 16 24 2e-4
Large 770M 1280 20 36 2e-4
Figure 1: Validation loss curves for communication interval $\tau=12$.
@@ 3 Convergence analysis
Randomized sign operator. In our analysis, we approximate the sign operator with a randomized analog. For any vector ${\bm{v}}=[v_{1},\ldots,v_{d}]^{\top}\in\mathbb{R}^{d}$ such that $\|{\bm{v}}\|\leq B$, we use the randomized sign operator $\mathcal{S}_{r}({\bm{v}})$, where the $j$-th component is defined as:
@@
We use these randomized sign operators as a linear and continuous analog of the original sign operator in expectation. Let $\mathbb{E}_{\mathcal{S}}$ denote the expectation taken over the randomness of the randomized sign operator, we have the following property.
@@
Then, by approximating sign with randomized operator $\mathcal{S}_{r}$ and using parameters $\gamma,\eta,\lambda,\beta_{1},\beta_{2}$ and $T$ as in Theorem 1, we have $\frac{1}{\tau T}\sum_{t=0}^{T-1}\sum_{k=0}^{\tau-1}\mathbb{E}[\|\nabla f(\bm{x}_{t,k})\|^{2}]=O(1/\sqrt{\tau T})$.
@@
We compare the instance in Theorem 2 with a similar algorithm, Federated Majority Vote signSGD with Stochastic Simple Momentum (Federated MV-sto-signSGD-SIM), analyzed by Sun et al. (2023). First, our algorithm uses aggregated local differences $(\bm{x}_{t,0}-\bm{x}_{t,\tau})$ for a global momentum update while MV-sto-signSGD-SIM uses local stochastic gradients to update local momentum before applying sign based majority vote. Second, Sun et al. (2023) prove its convergence to an $O(dR)$ neighborhood under the same Assumption 3, while Theorem 2, by using a simplified randomized sign $\mathcal{S}_{r}$, shows convergences to 0 with rate $O(1/\sqrt{T})$.

### v2

@@ Abstract
Pre-training Transformer models is resource-intensive, and recent studies have shown that sign momentum is an efficient technique for training large-scale deep learning models, particularly Transformers. However, its application in distributed training remains underexplored. This paper investigates a novel communication-efficient distributed sign momentum method with multiple local steps, to cope with the scenarios where communicating at every step is prohibitive. Our proposed method allows for a broad class of base optimizers for local steps, and uses sign momentum in the global step, where momentum is generated from differences accumulated during local steps. For generic base optimizers, by approximating the sign operator with a randomized version that acts as a continuous analog in expectation, we present a general convergence analysis, which specializes to an $O(1/\sqrt{T})$ rate for a particular instance. When local step is stochastic gradient descent, we show an optimal $O(1/T^{1/4})$ rate in terms of $\ell_{1}$ gradient norm for nonconvex smooth cost functions. We extensively evaluate our method on the pre-training of various sized GPT-2 models from scratch, and the empirical results show significant improvement compared to other distributed methods with multiple local steps.
@@ 1 Introduction
$\mathop{\mathrm{minimize}}_{\bm{x}\in\mathbb{R}^{d}}f(\bm{x}):=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{\bm{\xi}^{(i)}\sim\mathcal{D}_{i}}f_{i}(\bm{x},\bm{\xi}^{(i)}),$ (1)
where $f_{i}$ represents the cost function of a supervised learning model, $\bm{x}$ denotes the model parameters, and $\bm{\xi}^{(i)}$ refers to the data samples drawn from data distribution $\mathcal{D}_{i}$ on worker $i\in[n]$. In distributed training tasks, $\mathcal{D}_{i}$ can be a finite collection of samples allocated on worker $i$.
@@
The problem formulation (1) serves as a natural abstraction for distributed training (Lian et al., 2017; McMahan et al., 2017). A common approach to solving (1) involves workers first computing mini-batch gradients of $f_{i}$, followed by an all-reduce step to aggregate the local mini-batch gradients and obtain the global average. A global optimization step, such as Stochastic Gradient Descent (SGD) (Robbins and Monro, 1951), Momentum (Polyak, 1964), Nesterov’s Accelerated Gradient (NAG) (Nesterov, 1983), Adam (Kingma, 2014) and its weight decay variant AdamW (Loshchilov, 2017), or more recently Lion (Chen et al., 2024b), is then applied. This method is well suited for distributed learning scenarios where communication is inexpensive.
@@
However, in large-scale distributed training across GPU clusters, synchronized communication during the all-reduce step can become a bottleneck due to stragglers (Dean et al., 2012) or slow inter-node and inter-cluster communication, making per-step communication impractical. Several approaches have been proposed to mitigate this communication overhead. Decentralized methods, such as gossiping SGD (Jin et al., 2016) and consensus-based distributed SGD (Jiang et al., 2017), use approximate mini-batch gradient averaging instead of exact averaging to reduce idle time. Local step methods, like local SGD (Lin et al., 2020; Gorbunov et al., 2021b), local momentum (PR-SGD-Momentum) (Yu et al., 2019) and local Adam (Lu et al., 2023), reduce communication frequency to save on bandwidth. Additionally, some local step methods further optimize performance by overlapping communication and computation (Wang et al., 2019; Sun et al., 2024), utilizing stale synchronized model parameters for local steps to reduce the waiting time for synchronizing the latest local models. All these approaches, although reduce communication and run faster, add perturbations to the optimization process. These perturbations can lead to decreased performance in terms of training and validation losses under the same amount of computation (Wang et al., 2020). In this work, we focus on designing local step methods with excellent communication-performance tradeoffs.
@@
Several methods have been proposed to enhance the performance of local steps techniques. SCAFFOLD (Karimireddy et al., 2020) introduces a control variate that corrects the local SGD direction using global information, eliminating the dependency on local gradient dissimilarity in the convergence guarantees. One of the most relevant methods, BMUF (Chen and Huo, 2016), combines local steps with a global momentum step to accelerate distributed training, without the need for additional communication (which can also be a bottleneck in high-dimensional problem), such as transmitting the control variate in SCAFFOLD or the local momentum in PR-SGD-Momentum (Yu et al., 2019). SlowMo (Wang et al., 2019) extends BMUF by providing a general framework that operates on local base optimizers, such as local SGD and SGP (Assran et al., 2019), and periodically applies a global momentum step after running the local optimizer for several iterations.
@@
Beyond local step methods, momentum has long been recognized as a key technique for improving deep learning training, enhancing both optimization and generalization (Sutskever et al., 2013). More recently, Kunstner et al. (2023) suggest that sign momentum may be the key factor behind Adam’s significant advantage over SGD in training Transformers (Vaswani, 2017). Chen et al. (2024b) introduced the Lion (Evolved Sign Momentum) optimization method, discovered through program search, which outperforms the widely used Adam (Kingma, 2014) in various vision and autoregressive language modeling tasks. Furthermore, Kunstner et al. (2023), through experiments with different batch sizes, suggest that Adam’s significant performance advantage over SGD likely stems from its similarity to sign descent with momentum, particularly in Transformers used for language models.
@@
Inspired by the effectiveness of the sign momentum method in training large neural networks, particularly Transformers, and the success of global momentum steps in local step methods, a natural question is:
@@
Can we extend the sign momentum method to distributed training of Transformers with local steps?
@@ 1.1 Contributions
In this work, we answer this question affirmatively by proposing a framework for distributed sign momentum with local steps, supported by both empirical evaluations and theoretical justifications. Our contributions are as follows:
@@
    We introduce a new framework for distributed sign momentum with local steps. This framework is versatile, allowing any proper base optimizers for local steps, while using differences accumulated from these updates to adjust the momentum in the global sign momentum step. Details are provided in Algorithm 1.
@@
    For generic base optimizer, we provide a general theoretical analysis of the proposed framework by approximating the sign operator with a randomized continuous analog in expectation. When specializing the base optimizer as SGD, we obtain an in-expectation convergence rate of $O\big{(}1/\sqrt{T}\big{)}$ in terms of squared Euclidean norms, where $T$ is the total global steps.
@@
    When the base optimizer is SGD and the actual sign operator is used, we show an optimal in-expectation rate of $O(1/T^{1/4})$ measured in $\ell_{1}$ gradient norms, matching the best-known rate of centralized sign momentum method.
  4. (d)
    We extensively evaluate the performance of Algorithm 1 in pre-training GPT-2 models of various sizes from scratch, demonstrating consistent improvements over other distributed methods with multiple local steps and highlighting its strength in large-scale scenarios where communicating at every step is prohibitively costly.
@@ 1.2 Related work
Communication compression. In distributed optimization, there has been extensive research on compressing communication between workers to reduce overhead. This includes methods such as 1-bit sign SGD (Seide et al., 2014; Bernstein et al., 2018), gradient difference compression (Gorbunov et al., 2021a), compression with error-feedback (Karimireddy et al., 2019), and compression with error reset (Xie et al., 2020), among others. Although this work focuses on local step methods without communication compression, several 1-bit compression (sign-based) methods, such as SIGNUM (Bernstein et al., 2019) and Major Vote signSGD (Sun et al., 2023), apply (random) sign operations to local momentum and use majority voting for local signs. In contrast, Algorithm 1 employs full-precision communication during the global synchronization step, with the sign taken after aggregation.
@@
$\bm{x}_{t+1}=\bm{x}_{t}-\eta_{t}\text{sign}(\nabla f(\bm{x}_{t},\bm{\xi}_{t})),$ (2)
@@
$\begin{split}\bm{m}_{t+1}=\beta\bm{m}_{t}+(1-\beta)\nabla f(\bm{x}_{t},\bm{\xi}_{t}),\\ \bm{x}_{t+1}=\bm{x}_{t}-\eta_{t}\text{sign}(\bm{m}_{t+1}),\end{split}$ (3)
@@
Distributed sign momentum. The idea of distributed sign momentum is also studied in works including Liu et al. (2024a) and Tang and Chang (2024). Tang and Chang (2024) uses Lion for multiple local steps and average model and momentum periodically, and Liu et al. (2024a) computes the momentum term in Lion locally for once and uses global aggregation for model updates. Conversely, our proposed framework supports multiple local steps using any suitable base optimizer to generate local differences, which are then leveraged for sign momentum in the global step.
@@ 2 Distributed sign momentum
where $\bm{\xi}_{t,k}^{(i)}$ is a sample drawn from $\mathcal{D}_{i}$. As in Wang et al. (2019), base optimizers can also involve communication steps, such as SGP (Assran et al., 2019), and momentum steps. We denote the final update direction applied to worker $i$ at iteration $(t,k)$ as ${\bm{d}}_{t,k}^{(i)}$.
@@
Note that (6)-(8) mimics the update rule of Lion (Chen et al., 2024a) (also described in Section A in Appendix), by treating the differences accumulated from local steps as pseudo-stochastic gradients. In both (6) and (8), we scale $(\bm{x}_{t,0}-\bm{x}_{t,\tau})$ by $\frac{1}{\gamma_{t}}$ to make the momentum buffer independent of the learning rate $\gamma_{t}$, which can vary over time when using a learning rate schedule.
@@
9:     Update global parameter: $\qquad\bm{u}_{t+1}\leftarrow\beta_{1}\bm{m}_{t}+\frac{1-\beta_{1}}{\gamma_{t}}(\bm{x}_{t,0}-\bm{x}_{t,\tau})$ $\qquad\bm{x}_{t+1,0}\leftarrow\bm{x}_{t,0}-\eta\gamma_{t}\big{(}\text{sign}(\bm{u}_{t+1})+\lambda\bm{x}_{t,0}\big{)}$
@@
Distributed optimizer. We note that Algorithm 1 can be seamlessly integrated with the ZERO-series distributed optimizer (Rajbhandari et al., 2020) to handle large models. Consider a layered distributed optimization strategy: GPUs on a single node can be treated as a joint worker, utilizing a distributed optimizer like ZERO2 for local steps of the base optimizer, while the global buffers $\bm{x}_{t,0}$ and $\bm{m}_{t,0}$ are distributed across nodes. This approach allows more frequent communication for local steps to leverage faster intra-node communication, such as NVLinks.
@@ 3 Convergence analysis
3.1 Randomized sign operator
For a generic base optimizer that generates update directions satisfying Assumption 2 and 3, we present an general analysis that approximates the sign operator with a randomized analog. For any vector ${\bm{v}}=[v_{1},\ldots,v_{d}]^{\top}\in\mathbb{R}^{d}$ such that $\|{\bm{v}}\|\leq B$, we use the randomized sign operator $\mathcal{S}_{r}({\bm{v}})$, where the $j$-th component is defined as:
@@ 3.1 Randomized sign operator
We use these randomized sign operators as a linear and continuous analog of the original sign operator in expectation as in Safaryan and Richtárik (2021). Let $\mathbb{E}_{\mathcal{S}}$ denote the expectation taken over the randomness of the randomized sign operator, we have the following property.
@@
Table 1: Model configurations and peak learning rates
Size Embedding dimension # of heads Depth Peak LR
Small 125M 768 12 12 5e-4
Medium 355M 1024 16 24 2e-4
Large 770M 1280 20 36 2e-4
@@
Then, by approximating sign with randomized operator $\mathcal{S}_{r}$ and using parameters $\gamma,\eta,\lambda,\beta_{1},\beta_{2}$ and $T$ as in Theorem 1, we have $\frac{1}{\tau T}\sum_{t=0}^{T-1}\sum_{k=0}^{\tau-1}\mathbb{E}[\|\nabla f(\bm{x}_{t,k})\|^{2}]=O(1/\sqrt{T})$, specifically, it
$\quad+\frac{3n\tau^{2}L^{2}R^{2}(\sigma^{2}+3\tau\delta^{2})}{\eta^{2}T}+\sigma^{2}L\sqrt{\frac{n}{\tau T}}$
$\quad+\frac{4nL^{2}\sigma^{2}}{T}\Big{[}\big{(}\frac{\tau R}{\eta}-1\big{)}^{2}+\frac{2\beta^{2}}{1-\beta^{2}}\Big{]}+dLR^{2}\sqrt{\frac{n\tau}{T}}.$
@@
We compare the instance in Theorem 2 with a related algorithm, Federated Majority Vote signSGD with Stochastic Simple Momentum (Federated MV-sto-signSGD-SIM), analyzed by Sun et al. (2023). Federated MV-sto-signSGD-SIM incorporates similar algorithmic concepts but remains distinct from Algorithm 1. First, our algorithm uses aggregated local differences $(\bm{x}_{t,0}-\bm{x}_{t,\tau})$ for a global momentum update while MV-sto-signSGD-SIM uses local stochastic gradients to update local momentum. Second, Sun et al. (2023) uses randomized sign to compress worker to server communications, i.e., sign based majority vote, while we use full-precision communication among workers.
3.2 Sign operator
When the base optimizer is SGD, we also present its convergence analysis using actual sign for Algorithm 1.
Theorem 3.
Let Assumptions 1, 3 hold. In addition, the local update on worker $i$ satisfies that ${\bm{d}}_{t,k}^{(i)}=\nabla f_{i}(\bm{x}_{t,k}^{(i)},\bm{\xi}_{t,k}^{(i)})$, and $\forall\bm{x}\in\mathbb{R}^{d},\mathbb{E}_{\bm{\xi}^{(i)}\sim\mathcal{D}_{i}}[\|\nabla f_{i}(\bm{x},\bm{\xi}^{(i)})-\nabla f_{i}(\bm{x})\|^{2}]\leq\sigma^{2}$ for some $\sigma>0$. Taking $\eta=\frac{1}{LT^{3/4}}$, and $1-\beta=\frac{1}{\sqrt{T}}$, it holds that
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\big{[}\|\nabla f(\bm{x}_{t,0})\ \|_{1}\big{]}\leq\frac{L(f(\bm{x}_{0,0})-f_{*})}{\gamma T^{1/4}}$
$\quad+\frac{2\sqrt{d}}{T^{1/2}}\|\nabla f(\bm{x}_{0,0})\|+\frac{2d\gamma}{T^{1/4}}$
$\quad+\frac{2\sigma}{T^{1/4}}\sqrt{\frac{d}{\tau n}}+\frac{\sqrt{d}\tau R+\gamma d/2}{T^{3/4}}.$
Remark 2.
The convergence rate above in terms of $\ell_{1}$ gradient norm matches the rate of centralized sign momentum method, namely, update in (3), as shown in Sun et al. (2023). Specifically, in one of the leading terms $\frac{2\sigma}{T^{1/4}}\sqrt{\frac{d}{\tau n}}$, the benefit of multiple local steps and linear speedup are demonstrated, and it suggests our algorithm is preferable in large noise regime. We note that Sun et al. (2023) also shows similar convergence rate of federated MV-sto-signSGD-SIM), but only to an $O(\frac{dR}{\sqrt{n}})$ neighborhood under the same Assumption 3, due to its use of randomized signs for worker to server communication.